Operations Report – February 15

Crew 221 Operations Report 15Feb2020

SOL: 13

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 6:16 pm)

Hours run: 14:16

From what time last night: 7:16pm

To what time this morning: 5:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 70%

Diesel Reading – 73%

Station Propane Reading – 45%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 75 gals

Static to Loft Pump used – yes

Water in GreenHab – no water left in the greenhab yesterday, tank being refilled at this moment.

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours: 142.8

Beginning charge: 100%

Ending charge:

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: no

Hours: (before EVA) : 131.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging)

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 137.2

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: deep cleaned for crew handover

Summary of GreenHab operations: deep cleaned for crew handover, tank being refilled

Summary of ScienceDome operations: deep cleaned for crew handover Summary of RAM operations: RAM cleaned

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report.

Questions, concerns and requests to Mission Support: Full cleaning of habitats carried out, food inventory carried out.

GreenHab Report – February 15th

Crew 221 GreenHab Report 15FEB2020

GreenHab Officer: Jessica Todd

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on. Door opened today from 10:30 am to 3:48 pm.

Average temperatures: 18.25°C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 16.4 gallons

Daily water usage for research and/or other purposes: 0 gallon

Water in Blue Tank: empty (using static tank), being refilled tomorrow

Time(s) of watering for crops: 2

Changes to crops: The cucumber plant has a new cucumber, ready for harvesting in a few days. Harvested all the spinach chard, rocket and spinach from the grey pot by the door and reseeded. Harvested all icicle radish and reseeded with fresh radish. Should sprout in the next 2-3 weeks.

Narrative: This is my last report as GreenHab officer for Crew 221. It’s been a wonderful experience being entrusted with the care of these plants for the last two weeks. I don’t consider myself much of a gardener, but I’ve developed a real love for watching plants grow and harvesting my own greens from a garden I’ve helped to keep alive in the harsh environment of Mars! In preparation for the arrival of crew 222, today was GreenHab clean up! The whole GreenHab was swept and cleaned, all plants watered, and the remnants of our mungbeans and corn experiments packed up. Shannon came by and we discussed some final harvesting, so for the next few hours I harvested a huge pot of greens and a planter of radishes, and replanted them all. I went through and relabeled a few of the pots, including the chamomile and mint I planted a few days ago (enjoy some lovely fresh brew tea Crew 222!). This evening it was time for my final watering of the GreenHab. I introduced Crew 222’s GreenHab Officer Bleuenn to the ways of the GreenHab and tried to impart as much Mark Watney knowledge as I could. I hope she finds the job as rewarding as I have! Good luck my little Martian botany friends, you’re in safe hands xx

Harvest: 60g chard, 12g rocket, 26g radish greens.

Support/supplies needed: NA

Other Notes: The bottom row of pots on the central shelves behind the benches have not sprouted in two weeks and I can’t find any evidence of roots/shoots/seeds. I cleaned out the pots, removed the labls, and left the soil there for replanting. The GreenHab tank was empty this morning, so on Shannon’s instructions we used water from the tap outside the GreenHab that is fed by the static tank. Crew 222 started refilling the GreenHab tank today so hopefully tomorrow they will start their sim with a full 300 gallon tank.

EVA Report – February 14th

Crew 221 EVA Report 14FEB-2020

EVA #14

Author: Rawan Alshammari (Health and Safety Officer)

Purpose of EVA: Taking out robot for further testing and deployment of the crew engineer’s antenna to see how both systems perform in the field. Looking for geological rocks if applicable as secondary purpose.

Start time: 10:55am

End time: 1:20 pm

Narrative: Today we wanted to further test our robot and antenna to see how they perform on different terrain. We were able to set up the antenna successfully, and MOUSIE was ready for deployment. We were headed to the Moons, to send MOUSIE to collect some new samples. The moons were a perfect site to re-test MOUSIE. MOUSIE was able to access rock samples using her little robotic arm again.

Destination: White Moon, Beige Moon.

Coordinates (use UTM NAD27 CONUS): 4254800/515700

Participants: Rawan Alshammari (Health and Safety Officer), Maria Grulich (Commander).

Road(s) and routes per MDRS Map: Drove along Cow Dung Road (0110) and parked to walk Brahe Highway (1572) and Brahe Highway (1575)

Mode of travel: Driving and Walking

Sol Summary – February 14th

Crew 221 Sol Summary Report 14Feb-2020

Sol:12

Summary Title: Preparation to fly to Earth

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Today is Valentine’s day on Earth and also the last day for us on Mars. Yesterday the team played “King Mao” , a card game from Crew GreenHab Officer J. Todd and the team loved a lot, it is about figuring out rules that only one person knows, that is “King Mao”!

The crew woke up for the last time on Mars and celebrated it with scrambled eggs prepared again from crew journalist H. Kalucha. They were delicious!

After a great breakfast crew commander M.Grulich and crew HSO R. Alshammari prepared for the last EVA on Mars to the Moons! The EVA preparation went well and also the whole EVA was a complete success. They installed the antenna again in the field on top of a little hill. The rover was used in two different locations collecting very interesting samples as the surface was white and underneath red. The team also used acid to see if there are some carbonate components but no luck.

After the EVA was completed, the team had lunch and prepared some outreach videos to be shown back in our countries as well as doing last soil testing and checking on the microbiology samples.

On our last evening on Mars we will have a Valentine’s Day gift exchange as we draw secretly names of the crew members and prepared a Valentines gift! The crew commander also prepared a small surprise for the crew… See you back on Earth! It was a great two weeks and we achieved almost all our objectives and Mars treated us well! Until the next time… Ad Astra!

Anomalies in work: None

Update mouse in hab: no new mouse caught

Weather: sunny light winds

Crew Physical Status: Crew feeling good

Reports to be filed:

Sol Summary

Photo Report

Operations Report

Journalist Report

GreenHab Report

EVA Report 14

Support Requested: None

Commander Report – February 15th

Crew 221 Commander Report 15FEB2020

Author: Maria Grulich, Commander

Title:Returning to Mars

I can almost not believe it that my second rotation at MDRS is over. The first time I was part of the international emerging space leaders crew and I returned as the crew commander of an international crew from five different countries (Costa Rica, Australia, Canada, Germany, Kuwait) where not everyone knew each other. Arriving back at the station it was as I never left, I immediately felt familiar and back at home as if my friends from crew 205 are still around, especially since crew 205 left us some surprise cookies in the freezer! The first week of MDRS we were very productive going on a lot of EVAs and I was surprised that the weather stayed stable for the two weeks we were here!

The second week passed by very quickly as we realised that two weeks is not so much time when you still have to finish so many projects while keeping a Mars Station up and running. The projects included growing black corn and sprouting mung beans, investigating microbes in Martian soil, testing salinity of the soil and building a robot to assist during EVAs, while filling in surveys for human factor studies and following the international voice protocol used on the International Space Station in the hab as well as on EVAs.

Being the commander of an international crew is very enriching, in particular, as we organised cultural night every other day where we learned how to speak Australian, danced to “Nutbush City Limits”, dressws as Kuwaities, got a history lesson about India and Canada, ate nice food from Costa Rica, and danced to German Octoberfest songs and learned how to make the longest words in the World in German. We managed to entertain ourselves with games, sport and meditation sessions and a lot of good conversations, which always raised the question is it “cultural” or “religion”.

It was an honor to be the crew commander of Crew 221 Earth2Mars!

I want to thank all the crew members, Hemani Kalucha Crew Journalist alias “Juno”, Louis D. Monge Crew Engineer alias “Frankie”, Jessica Todd Crew Greenhab Officer alias “Goblin”, Rawan Alshammari Health and Safety Officer alias “Luna”, and Ghanim Alotaibi Crew Astronomer alias “Astro” for all their effort and being so great in achieving almost all our objectives! We came as strangers and we leave as friends! The team is already planning their reunion in Costa Rica in a year we really hope that this will happen in the meantime almost all our crew members are part of SGAC (Space Generation Advisory Council) and will meet on their events and space conferences for sure sharing now this amazing experiences of being two weeks on Mars together!

Until the next time… Ad Astra!

Crew Commander Grulich alias “Ewok” from Crew 221 Earth2Mars signing off….

MDRS Station

Sol Summary – February 15th

Crew 221 Sol Summary Report 15Feb-2020

Sol:13

Summary Title: Touch down Earth awaits

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol Summary: The team prepared the Mission summary of Crew 221 Earth2Mars, reporting about the mission and what we achieved from our original plan. Besides a few drawbacsk on the weather station and geology testing we fulfilled all our objectives.

After the work was done the team gathered together and went through our personal mission goals and our original expectations. We also asked everyone before the mission what are their special skills and after two weeks spending each minute together, they may have developed new once. After this the crew commander handed every member of the team a self-written certificate about achieving to live two weeks on Mars together at the Mars Desert Research Station and handing them the “New Horizons” Pin of the European mission of Alexander Gerst a German Astronaut as a reminder what they achieved. To finish up the Valentine’s Day we exchanged our gifts that we brought from our countries before to give to our secret Valentine! Even our seventh crew member in our heart who could not come send her gift and was very well received.

The crew went to bed the last time on Mars….

To wake up on Earth Saturday morning!

The first morning on Earth the crew commander, crew engineer, and crew Greenhab officer ran out with their pyjamas to breath fresh Earth air without an EVA suit!

Once the rest of the crew was awake, we had breakfast and get started to clean as we required all hands-on deck. After a few hours everything was cleaned and we were ready to meet crew 222 at the rock shop!

Crew 221 prepared some soup and one last time crew commander made some German bread for both crews.

Crew 221 went out for a walk to conquer the North Ridge as it has become a tradition started by crew 205 and continued by crew 219 and now 221.

The rest of the time was spent to train crew 222. The official handover was done at 4:30pm!

Both crews will be heading to Stan’s for hamburgers at 6:00pm before our last time sleeping in our crew quarters and leaving the station tomorrow at 8:00am!

Crew 221 signing off…

Anomalies in work: None

Update mouse in hab: no new mouse caught

Weather: morning cloudy light winds, afternoon partially sunny light winds

Crew Physical Status: Crew feeling good

Reports to be filed:

Sol Summary

Photo Report

Operations Report

Commander Report

GreenHab Report

Support Requested: None

Operations Report – February 14th

Crew 221 Operations Report 14Feb2020

SOL: 12

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 7:16 pm) 89%

Hours run: 10:40

From what time last night: 8:20pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 73%

Diesel Reading – 73%

Station Propane Reading – 47%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 145 gals

Static to Loft Pump used – yes

Water in GreenHab – 157.5 gals used and no water left in the greenhab at this moment, please refer to greenhab report.

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 142.8

Beginning charge: 100%

Ending charge: 47%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 131.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 53%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 137.2

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment finished, watering of the plants carried out as per procedure.

Summary of ScienceDome operations: one oven is operating in the lab, the incubator is non-functional, microbiology experiments going on and cultures are being grown in agar plates, all activities set to end by tomorrow.

Summary of RAM operations: RAM cleaned

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report.

Questions, concerns and requests to Mission Support: Cleaning of habitats, food inventory carried out.

Science Report – February 14th

Crew 221 Science Report 14-FEB-20
Crew Science Officer: Hemani Kalucha

1. Microbiology Research
Today, we tested all our samples for pH and salinity.

2. ISRU project
Corn was Harvested.

3. Mung Beans

Nothing Further to report.

4. Radio Communication

Nothing further to report.

5. Geology Tests

Our geology kit was only found today in Shannon’s trailer. We conducted preliminary HCl tests on some samples.